Brandon is 10 years older than sonya. in two years Brandon will be 3 times as old as sonya. how old are they now
Answer by rothauserc(4718)

let Sonya be x years old and Brandon is x+10 years old, then
x+10 +2 = 3*(x+2)
x + 12 = 3x +6
2x = 6
x = 3
Sony is 3 and Brandon is 13
**************************
check the answer
3+10 +2 = 3*(3+2)
15 = 3*5
15 = 15
our answer checks